# Emergence in graphs with near-extreme constraints

### v2

@@ 1.4. Definitions
A graphon is said to be k-podal if we can partition the interval $[0,1]$ into measurable sets $I_{1},\ldots,I_{k}$ such that $g(x,y)$ is constant on each “rectangle” $I_{i}\times I_{j}$. By picking an appropriate representative of the equivalence class of $g$, we can assume that each $I_{i}$ is an interval. We refer to the sets $I_{i}$ as podes. If $g$ is $k$-podal for some integer $k$, we say that $g$ is multipodal. We often use the words bipodal and tripodal to mean 2-podal and 3-podal. We say that an $(n+m)$-podal graphon has $(n,m)$ symmetry if $g$ is invariant under permutation of the first $n$ podes and is also invariant under permutation of the remaining $m$ podes. The space of such $(n,m)$-symmetric graphons is only 6-dimensional (or less if $n$ or $m$ is less than 2), with parameters specifying the relative size of the two kinds of podes and the value of $g(x,y)$ on the five (or fewer) equivalence classes of rectangles $I_{i}\times I_{j}$. A graphon with $(2,0)$ symmetry is said to be symmetric bipodal.
In a model with constraints on the densities $\varepsilon,\eta$ of edges and another graph $H$, the phase diagram is the set of all achievable values $(e,t)$ of $(\varepsilon,\eta)$. A phase is a connected open set in the phase diagram where the optimal graphon is unique and is a real analytic function of $(e,t)$. More precisely, for each finite graph $K$ (such as a square or tetrahedron or pentagon), there is a function $\kappa(g)$, similar to $\varepsilon(g)$ and $\tau(g)$, that gives the density of $K$’s in $g$. Within a phase, each such functional applied to the optimal graphon is required to yield an analytic function of $(e,t)$. In practice, analytic dependence of the optimal graphon $g$ on $(e,t)$ is usually proven by describing $g$ (and thus the density of every subgraph $K$) via a finite set of parameters and then showing that these parameters are analytic functions of $(e,t)$. A phase transition occurs where the density of some $K$ is not analytic or is not even defined, such as when the optimal graphon is not unique. Phase transitions have only been shown to occur on boundaries of phases.
@@ 1.5. Detailed results
The following theorems give a simple description of what happens near almost all points along the boundary of the Razborov triangle. We prove that the unique optimal graphons in the phases near the boundary are multipodal. The cited theorems in the text include additional estimates on how the podes of the optimal graphons scale as the constraints approach the boundary.
@@
For each fixed $e<1/2$ and all $t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t)$ is unique and is symmetric bipodal, with parameters that vary analytically with $(e,t)$.
@@
Let $n\geq 1$ be an integer. For every $e\in\left(\frac{n}{n+1},\frac{n+1}{n+2}\right)$, with corresponding minimal triangle density $t_{0}$ (depending on $e$), and for all $\Delta t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t_{0}+\Delta t)$ is unique and $n+2$-podal, with $(n,2)$ symmetry and with parameters that vary analytically with $(e,t)$.
@@
All of the phases above the scallops proven in Theorems 1 and 2 have unique optimal graphons with distinct symmetries and cannot be analytically continued to one another.
@@
All of these theorems can be viewed as extensions of extremal graph theory. Pikhurko and Razborov’s results [35] determine unique optimal graphons on the boundary of Figure 1 [37]. Theorems 14 describe their unique extensions nearby. This is in contrast to our earlier results that relied on perturbing Erdős-Rényi graphs, or on perturbing known results on the $e=1/2$ line for $t<e^{3}$, all well within the interior of the Razborov triangle.
2. The semi-local “worth” functional
@@
In this section we develop the concept of “worth”. This is a quantity associated with columns of a graphon. We will see that all columns of an optimal graphon must maximize worth. Working with columns, rather than just with the value $g(x,y)$ of the graphon at each point separately, gives us the analytic control needed to prove our main theorems.
We recall the basic theory of Lagrange multipliers. Let $g$ be a graphon that maximizes the Shannon entropy $S(g)$ for particular values $(e,t)$ of $(\varepsilon(g),\tau(g))$. We call an infinitesimal change to the graphon reversible if it is possible to make the opposite change, for which the infinitesimal changes $dS$, $d\varepsilon$, and $d\tau$ are minus those associated with the original change. For instance, increasing the value of a graphon in a portion of $[0,1]^{2}$ from 0.5 to slightly above 0.5 is reversible, since we could just as well lower the value slightly below 0.5. Likewise, expanding the area of the region where $g(x,y)=0.5$ (at the expense of a region where $g(x,y)$ takes on a different value) is reversible, since we could just as well shrink the region. However, increasing the value of a graphon from 0 to a positive number, or increasing the area of a region from 0 to a positive number, are irreversible changes.
Note that the terms reversible and irreversible do not refer to changes in time, as we are not working with any notion of dynamics. Rather, we can visualize infinitesimal changes as tangent vectors to the space of possible graphons. In this heuristic picture, irreversible changes correspond to inward-pointing vectors at the boundary of the space, while reversible changes either refer to points in the interior or are tangent to the boundary.
When considering reversible changes to an optimal graphon, the infinitesimal changes $dS$, $d\varepsilon$ and $d\tau$ to the entropy, edge density and triangle density cannot be linearly independent at $g$. If they were, we could find a reversible change that increased $S$ while leaving $\varepsilon$ and $\tau$ fixed, contradicting the optimality of $g$. There must therefore exist constants $\alpha$ and $\beta$ (depending on $g$) such that, for any reversible infinitesimal change to $g$,
(5) $dS=\alpha d\varepsilon+\frac{\beta}{3}d\tau.$
Equation (5) applies not only at points where $d\varepsilon$ and $d\tau$ are linearly independent, but also on the Erdős-Rényi curve, where $d\varepsilon$ and $d\tau$ are collinear. However, at such points the constants $(\alpha,\beta)$ are not uniquely defined. Rather, each Erdős-Rényi graphon is associated with infinitely many possible values of $(\alpha,\beta)$.
For irreversible changes, the equality is replaced by an inequality: $dS\leq\alpha d\varepsilon+\frac{\beta}{3}d\tau.$ There is no problem with such moves costing us entropy relative to the prevailing exchange rate for $\varepsilon$ and $\tau$, since the opposite move is not possible, but they cannot gain entropy.
This principle is sometimes described in terms of a functional
@@
To first order, reversible changes to $g$ cannot change $F$, while irreversible changes can decrease $F$ but cannot increase $F$.
The functionals $S$ and $\varepsilon$ are local, in that there is a contribution from each point $(x,y)\in[0,1]^{2}$ and we integrate the local contributions to get the global quantity. If the triangle density were also local, then $F$ would be the integral of a local density and our variational equations would come from setting the derivative of this density with respect to $g(x,y)$ equal to zero. That is the typical situation when doing calculus of variations, especially in classical and quantum field theory, with the global action being the integral of a local Lagrangian density [5].
@@
is not local. Rather, it involves interactions between the values of $g$ at three different points. To accommodate this complication, we consider (reversible and irreversible) changes to entire columns of $g$ and not just to isolated points. We work up to that in stages, first considering infinitesimal changes to $g$ near a single point, then considering macroscopic changes to $g$ near that point, and finally considering macroscopic changes on an entire column.
If we vary slightly the value of $g(x,y)$ in a neighborhood of a point, then standard methods apply. The functional derivatives of $S$, $\varepsilon$ and $\tau$ with respect to $g(x,y)$ are $H^{\prime}(g(x,y))$, $1$ and $3G(x,y)$, respectively, where
[…]
We conclude that
(9) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y)$
at every point where $g(x,y)$ is continuous. As for points of discontinuity, equation (9) must be understood in a weak sense. If we multiply $H^{\prime}(g(x,y))$ and $\alpha+\beta G(x,y)$ by an arbitrary test function and integrate, the two integrals must be the same.
We can also consider changing the value of $g(x,y)$ by a larger amount, albeit on an infinitesimally small region $R$. Note that this is still an infinitesimal change, in that the resulting (infinitesimal) changes in $\varepsilon$ and $\tau$ can be turned into changes in $S$ by making infinitesimal changes to $g$ elsewhere. The change in $F$ is proportional to the area of $R$ times the change in
(10) $V(x,y)=H(g(x,y))-\alpha g(x,y)-\beta G(x,y)g(x,y).$
The derivative of $V(x,y)$ with respect to $g(x,y)$ is precisely
(11) $H^{\prime}(g(x,y))-\alpha-\beta G(x,y),$
so maxima of $V$ are automatically solutions to (9). Furthermore, the second derivative of $V$ with respect to $g(x,y)$ is always negative, so solutions to (9) always maximize $V$.
An optimal graphon must maximize $V$ at each point, but that local optimization is not sufficient. To account for the non-locality of the $\tau$ function, we define the worth of a function $C:[0,1]\to[0,1]$ to be
[…]
@@
If a graphon $g$ maximizes the Shannon entropy $S(g)$ for given values of edge and triangle density, then every column of $g$ maximizes $W$, using the values of $\alpha$ and $\beta$ determined by equation (5) (for instance via equation (9)). In particular, all columns of $g$ must have the same worth.
@@
[…]
Using Theorem 5, we can now prove Theorems 1, 2 and 4.

### v3

@@ 1.3.4. Multipodality, phases and phase transitions
A graphon is said to be k-podal if we can partition $[0,1]$ into $k$ measurable sets $I_{1},\ldots,I_{k}$ such that $g(x,y)$ is constant on each “rectangle” $I_{i}\times I_{j}$. We refer to the sets $I_{i}$ as podes. If $g$ is $k$-podal for some integer $k$, we say that $g$ is multipodal. We often use the words bipodal and tripodal to mean 2-podal and 3-podal. We say that an $(n+m)$-podal graphon has $(n,m)$ symmetry if $g$ is invariant under permutation of $n$ of the podes and is also invariant under permutation of the remaining $m$ podes. A graphon with $(2,0)$ symmetry is said to be symmetric bipodal.
A phase is a connected open set in the interior of the Razborov triangle (Figure 1) in which the reduced optimal graphon is unique and is a real analytic function of $(e,t)$ in the following sense. Within a phase, and for each simple graph $F$, $\tau_{F}$ of the optimal graphon is required to be an analytic function of $(e,t)$. In practice, the analyticity of all densities $\tau_{F}$ is proven by first showing that the optimal graphon is unique and multipodal of a certain form and then showing that the finitely many parameters needed to describe this multipodal graphon are analytic functions of $(e,t)$.
A phase transition occurs where some $\tau_{F}$ is not analytic or is not even defined, such as where the optimal graphon is not unique. Phase transitions have only been shown to occur on boundaries of phases, on curves.
@@ 1.4. Detailed results
The following theorems give a simple description of what happens near almost all points along the boundary of the Razborov triangle, Figures 1 and 2. We prove that the unique optimal graphons in the phases near the boundary are multipodal. The cited theorems in later sections include additional estimates on how the parameters of the optimal graphons scale as the constraints approach the boundary.
@@
For each fixed $e<1/2$ and all $t$ sufficiently small, the optimal (reduced) graphon in $\widetilde{\mathcal{W}}_{e,t}$ is unique and is symmetric bipodal, with parameters that vary analytically with $(e,t)$.
@@
Let $n\geq 1$ be an integer. For every $e\in\left(\frac{n}{n+1},\frac{n+1}{n+2}\right)$, with corresponding minimal triangle density $t_{0}$ (depending on $e$), and for all $\Delta t$ sufficiently small, the optimal (reduced) graphon in $\widetilde{\mathcal{W}}_{e,t_{0}+\Delta t}$ is unique and $n+2$-podal, with $(n,2)$ symmetry and with parameters that vary analytically with $(e,t)$.
@@
All of the phases above the scallops proven in Theorems 1 and 2 have unique optimal reduced graphons with distinct symmetries and cannot be analytically continued to one another.
@@
All of these theorems can be viewed as extensions of extremal graph theory. Pikhurko and Razborov’s results [40] determine unique, entropy-optimal graphons on the boundary of Figure 1 [42]. Theorems 14 describe the infinite number of distinct neighboring phases.
2. Lagrange multipliers and the “worth” functional
@@
In this section we develop the concept of “worth”. This is a quantity associated with columns of a graphon, that is with the functions $g_{x}(y)=g(x,y)$ of one variable defined by fixing $x$ and allowing $y$ to float. We will see in Theorem 11 that the columns of an entropy optimal graphon must maximize worth. Working with columns, rather than just with the value $g(x,y)$ of the graphon $g$ at each point $(x,y)$ separately, gives us the analytical control needed to prove our main theorems.
To accomplish this we need a theory of Lagrange multipliers. For these purposes, we treat graphons as elements of $L^{2}([0,1]^{2})$ with the $||\cdot||_{2}$-norm. The $L^{2}$-topology is finer than the topology defined by the cut distance. All $L^{2}$-limits are limits in the cut metric, but not vice-versa. In order to combine our machinery with established theorems about reduced graphons, we will eventually need to prove that certain sequences converge in $L^{2}$ and not just in the cut distance.
We develop Lagrange multipliers in several steps. First we develop the theory for a certain class of variations where the standard theory of functional derivatives applies. This determines our Lagrange multipliers $(\alpha,\beta)$. We then show that, for other $L^{2}$-small changes to an optimal graphon, a certain functional involving $\alpha$ and $\beta$ cannot increase to leading order in the size of the change. We compute the effect of changing a small set of columns of our graphon and express the difference in this functional in terms of the worths of the old and new columns. Since the functional cannot increase, we conclude that the columns of the original (optimal) graphon must all maximize worth.
The simplest changes involve varying the value of our graphon $g$ gradually at each point, as indicated by a bounded symmetric function $g_{1}:[0,1]^{2}\to\mathbb{R}$. We consider a family of graphons $g_{s}$ of the form
(9) $g_{s}(x,y)=g_{0}(x,y)+sg_{1}(x,y).$
To remain in $\mathcal{W}$ we must have $0\leq g_{0}(x,y)+sg_{1}(x,y)\leq 1$ for all sufficiently small $s$. This can always be achieved by choosing $g_{1}$ to be supported on a subset of $0,1]^{2}$ on which $g_{0}(x,y)$ is bounded away from 0 and 1. For now, we do not consider variations that change $g(x,y)$ at points where $g_{0}(x,y)=0$ or 1.
The resulting changes to the edge density, triangle density and Shannon entropy are:
(10) $\Delta\varepsilon:=\varepsilon(g_{s})-\varepsilon(g_{0})$ $=$ $s\iint g_{1}(x,y)\,dx\,dy+o(s),$
(11) $\Delta\tau:=\tau(g_{s})-\tau(g_{0})$ $=$ $s\iint 3G(x,y)g_{1}(x,y)\,dx\,dy+o(s),$
(12) $\Delta S:=S(g_{s})-S(g_{0})$ $=$ $s\iint H^{\prime}(g_{0}(x,y))g_{1}(x,y)\,dx\,dy+o(s),$
where
(13) $G(x,y)=\int_{0}^{1}g_{0}(x,z)g_{0}(y,z)\,dz,$
and $H^{\prime}(u)=\ln\left(\frac{1-u}{u}\right)$ is the derivative of $H$ defined in equation (8). The quantities $H^{\prime}(g_{0}(x,y))$, 11 and $3G(x,y)$ are called the functional derivatives of $S(g)$, $\varepsilon(g)$, and $\tau(g)$ with respect to $g$ at $g=g_{0}$ and are sometimes denoted
(14) $\frac{\delta S(g)}{\delta g(x,y)},\qquad\frac{\delta\varepsilon(g)}{\delta g(x,y)},\qquad\hbox{and}\qquad\frac{\delta\tau(g)}{\delta g(x,y)}.$
These functional derivatives are elements of $L^{2}(A)$, where $A$ is the subset of $0,1]^{2}$ where $0<g(x,y)<1$, and are only defined up to sets of measure zero.
Remark 5.
The graphon $g$ is an element of $L^{2}([0,1]^{2})$ and so is not literally a function. Rather, it is an equivalence class of functions that agree off sets of measure zero. Similarly, the “function” $G(x,y)$, which gives the inner product of the columns $g_{x}$ and $g_{y}$, is only defined up to sets of measure zero. If $\bar{g}:[0,1]^{2}\to[0,1]$ is a symmetric function that agrees with $g$ apart from a negligible set, then Fubini’s Theorem says that, for almost every $x$, the columns $g_{x}$ and $\bar{g}_{x}$ agree except on a negligible subset of $[0,1]$, and in particular represent the same “function” in $L^{2}([0,1])$. So for almost every pair $(x,y)\in[0,1]^{2}$, the inner product of $g_{x}$ and $g_{y}$ is the same as the inner product of $\bar{g}_{x}$ and $\bar{g}_{y}$. That is, the functions $G(x,y)$ computed from $g$ and $\bar{g}$ agree except on a set of measure zero.
Lemma 6.
If $g$ is not a constant graphon, then $G(x,y)$ is not constant.
Proof.
By Cauchy-Schwarz,
(15) $G(x,y)\leq\sqrt{G(x,x)G(y,y)}\leq\max(G(x,x),G(y,y)).$
Proving by contradiction, the only way for $G(x,y)$, $G(x,x)$ and $G(y,y)$ to all be equal is if the columns of $g$ at $x$ and $y$ are identical. But if all the columns of $g$ are identical (and likewise all of the rows, since $g$ is symmetric), then $g$ is a constant graphon. ∎
Theorem 7.
Let $g$ be an entropy-maximizing graphon, subject to the constraints $\varepsilon(g)=e$ and $\tau(g)=t$. If the function $G(x,y)$ is not constant on the set of points $(x,y)$ where $0<g(x,y)<1$, then there exist unique Lagrange multipliers $\alpha$ and $\beta$ such that
(16) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y)$
almost everywhere. Furthermore, the function $g(x,y)$ is bounded away from 0 and 1.
Proof.
Since $G(x,y)$ is not constant on the set $0<g(x,y)<1$, $G(x,y)$ is not constant on the set $\epsilon<g(x,y)<1-\epsilon$ for all sufficiently small $\epsilon$, so the functional derivatives of $\varepsilon$ and $\tau$ are linearly independent on this set, so we can vary $\varepsilon$ and $\tau$ independently by choosing appropriate functions $g_{1}(x,y)$ that are supported on this set. Similarly, if the three functions $\frac{\delta\varepsilon}{\delta g}$, $\frac{\delta\tau(g)}{\delta g(x,y)}$, and $\frac{\delta S(g)}{\delta g(x,y)}$ were linearly independent on the set of points where $0<g(x,y)<1$, then we could vary $\varepsilon$, $\tau$, and $S$ independently, and in particular we could increase $S$ while keeping $e$ and $t$ exactly constant. Since this contradicts the optimality of $g$, we conclude that $\frac{\delta S(g)}{\delta g(x,y)}$ is a linear combination of $\frac{\delta\varepsilon}{\delta g}$ and $\frac{\delta\tau(g)}{\delta g(x,y)}$, so we can write
(17) $\frac{\delta S}{\delta g(x,y)}=\alpha\frac{\delta\varepsilon}{\delta g(x,y)}+\frac{\beta}{3}\frac{\delta\tau}{\delta g(x,y)}.$
Since 1 and $G(x,y)$ are linearly independent, the coefficients $\alpha$ and $\beta$ are uniquely defined. Plugging the functional derivatives into equation (17) then gives (16), which applies at points where $0<g(x,y)<1$.
Since $\alpha$ and $\beta$ are finite and since $G(x,y)$ is bounded, $H^{\prime}(g(x,y))$ is bounded, so $g(x,y)$ cannot be arbitrarily close to 0 or 1. Either $g(x,y)$ equals 0 or 1 (in which case equation (16) does not apply) or $g(x,y)$ is bounded away from 0 or 1.
We now eliminate the first possibility. If $g(x,y)=0$ or 1 on a set of positive measure, we can change the value of $g(x,y)$ at such points away from 0 and 1, thereby increasing $S$ greatly while only changing $\varepsilon(g)$ and $\tau(g)$ slightly. Since we can vary $\varepsilon$ and $\tau$ independently with an appropriate choice of $g_{1}$, we can restore the original values of $\varepsilon$ and $\tau$ with the resulting change in $S$ being governed by (17). The total effect is to increase $S$ while leaving $\varepsilon$ and $\tau$ fixed, which is a contradiction. ∎
The crux of the proof is the same as the derivation of Lagrange multipliers in finite dimensions. The fact that linearly independent derivatives give us the ability to vary $\varepsilon$ and $\tau$ (and possibly $S$) independently is just the inverse function theorem in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. The treatment of the points where $g(x,y)=0$ or 1 is essentially the same as what we do to maximize a function on the boundary of a domain in $\mathbb{R}^{n}$. Our setting is the infinite-dimensional space of graphons, but the core arguments are just finite-dimensional calculus.
Theorem 7 was stated in terms of functional derivatives, but we can also speak in terms of small changes to a graphon. As long as the functional derivatives of $\varepsilon$ and $\tau$ are linearly independent on the set of points where $0<g(x,y)<1$, we can vary $\varepsilon$ and $\tau$ independently by choosing an appropriate $g_{1}(x,y)$, with the resulting change in $S$ being given by
(18) $\Delta S=\alpha\Delta\varepsilon+\frac{\beta}{3}\Delta\tau+o(s\|g_{1}\|_{2}).$
Remark 8.
But what if a graphon $g$ fails to meet the assumptions of Theorem 7, with $g(x,y)$ equaling 0 or 1 on part of the unit square and $G(x,y)$ being constant everywhere else? That is, what if we cannot vary $\varepsilon$ and $\tau$ independently at $g$? We call such graphons singular entropy maximizers. These certainly appear on the boundary of the Razbarov triangle. In the interior they appear on the Erdős-Rényi curve $t=e^{3}$ and may appear elsewhere, so sometimes extra work is needed to exclude them. (See [26, Theorem 4.1] for such an argument at a specific phase transition.)
Fortunately for us, such singular graphons do not appear in any of the regions being studied in this paper. We postpone an explanation of this fact to the end of this section and return to the more typical situation of nonsingular entropy maximizers that satisfy the assumptions of Theorem 7.
We now consider infinitesimal changes to a graphon obtained by changing $g(x,y)$ by a “macroscopic amount” on a set of infinitesimal measure $s$. For instance, we might vary the sizes of the podes in a multipodal graphon. Such changes are not covered by Theorem 7. Nonetheless, they satisfy Lagrange-like inequalities.
Proposition 9.
Suppose that $g_{0}$ is a non-constant and non-singular entropy maximizer with Lagrange multipliers $\alpha$ and $\beta$. Let $g_{s}$ be a family of graphons obtained by changing $g_{0}$ on sets of measure $s$. Then
(19) $\Delta S\leq\alpha\Delta\varepsilon+\frac{\beta}{3}\Delta\tau+o(s),$
where $\Delta S=S(g_{s})-S(g_{0})$, $\Delta\varepsilon=\varepsilon(g_{s})-\varepsilon(g_{0})$, and $\Delta\tau=\tau(g_{s})-\tau(g_{0})$.
Proof.
Supposing this to be false, we will construct a variation of $g_{0}$ that has the same edge and triangle densities but more entropy. As previously noted, we can adjust $\varepsilon$ and $\tau$ independently by adding a function supported on $\epsilon<g(x,y)<1-\epsilon$, with the resulting changes in $S$ given by equation (18). Applying these changes to $g_{s}$ to restore the original values of $\varepsilon$ and $\tau$, we should get a change in $S$ that is at least $\Theta(s)-o(s)>0$, which is a contradiction.
To complete this argument, we must bound the cross terms from adding a function of pointwise size $O(s)$ to $g_{0}$ and in changing $g_{0}$ on a set of measure $s$. Changing $g(x,y)$ by $O(s)$ can only change $H(g(x,y))$ by $O(s\ln(s))$ and can only change $g(x,y)g(y,z)g(x,z)$ by $O(s)$. Integrating over a region of size $O(s)$, this can change the entropy by $O(s^{2}\ln(s))$, the edge density by $O(s^{2})$ and the triangle density by $O(s^{2})$, resulting in an $o(s)$ change to $S-\alpha\varepsilon-\frac{\beta}{3}\tau$, which can be absorbed into the $o(s)$ error term in equation (18). ∎
The estimate (19), combined with the theory of Lagrange multipliers for pointwise-small changes (Theorem 7 as summarized in equation (18)) can be described in terms of a functional
@@
If $g$ is a nonsingular constrained entropy maximizer, then pointwise small changes to $g$ cannot change $F$ to first order, while macroscopic changes on small sets can decrease $F$ but cannot increase $F$ to first order.
The functionals $S$ and $\varepsilon$ are local, in that there is a contribution from each point $(x,y)$ in $0,1]^{2}$ and we integrate the local contributions to get the global quantity. If the triangle density were also local, then $F$ would be the integral of a local density and our variational equations would come from setting the derivative of this density with respect to $g(x,y)$ equal to zero. That is the typical situation when doing calculus of variations, especially in classical and quantum field theory, with the global action being the integral of a local Lagrangian density [5].
@@
(21) $\tau(g)=\iiint g(x,y)g(y,z)g(z,x)\,dx\,dy\,dz$
is not local. It involves interactions between the values of $g$ at the three points $(x,y)$, $(x,z)$ and $(y,z)$. To accommodate this complication it is useful to consider macroscopic changes to entire columns. We define a quantity that measures the effect of such changes.
Definition of worth: Let $C$ be a possible column of a graphon $g$. That is, $C:[0,1]\to[0,1]$ is a Borel measurable function. The worth of $C$ is
(22) $W(C)=2\int_{0}^{1}H(C(y))\,dy-2\alpha\int_{0}^{1}C(y)\,dy-\beta\iint C(y)C(z)g(y,z)\,dy\,dz.$
Note that this depends explicitly on the graphon $g$ as well as on $C$ and the Lagrange multipliers $\alpha$ and $\beta$.
Proposition 10.
Let $(\alpha,\beta)$ be specified and suppose that $\tilde{g}(x,y)=g(x,y)$ except when $x$ or $y$ lies in a set $I$ of measure $s$. Then
(23) $F(g)-F(\tilde{g})=\int_{I}W(g_{x})-W(\tilde{g}_{x})\,dx+O(s^{2}),$
where $g_{x}$ and $\tilde{g}_{x}$ are columns of $g$ and $\tilde{g}$.
Proof.
The quantities $\varepsilon(g)$, $S(g)$ and $\tau(g)$ are all double or triple integrals over $0,1]^{2}$ or $0,1]^{3}$. The integrands for $g$ and $\tilde{g}$ are identical except where one of the variables lies in $I$. To compute $F(g)-F(g^{\prime})$, we must only keep the contributions of $x\in I$, multiply by 2 or 3 to allow for the similar contributions of $y\in I$ or $z\in I$, and make adjustments for where two or three variables are in $I$. Since the integrand is bounded and the set of points where multiple variables lie in $I$ only has measure $O(s^{2})$, we can compute F(g)F(g~)F(g)-F($\tilde${g}) to within $O(s^{2})$ by assuming that $x\in I$ and leaving $y$ and $z$ free.
We begin with the edge density:
(24) $\varepsilon(g)-\varepsilon(g^{\prime})=2\int_{x\in I}\int_{y\in[0,1]}g_{x}(y)-\tilde{g}_{x}(y)\,dy\,dx+O(s^{2}).$
The entropy is similar:
(25) $S(g)-S(\tilde{g})=2\int_{x\in I}\int_{y\in[0,1]}H(g_{x}(y))-H(\tilde{g}_{x}(y))\,dy\,dx+O(s^{2}).$
The triangle density has a prefactor of 3 instead of 2 because it is a triple integral:
@@
(27) $=$ $3\int_{x\in I}\iint_{y,z\in[0,1]}\left(g_{x}(y)g_{x}(z)-\tilde{g}_{x}(y)\tilde{g}_{x}(z)\right)g(y,z)\,dy\,dz\,dx+O(s^{2}),$
where in the last line we used the fact that $\tilde{g}(y,z)$ only differs from $g(y,z)$ when $y\in I$ or $z\in I$. Multiplying the change in $\varepsilon$ by $\alpha$ and the change in $\tau$ by $\beta/3$ and adding terms, we get (23). ∎
@@
If $g$ is a nonsingular entropy maximizer with Lagrange multipliers $(\alpha,\beta)$ then $g$ agrees off a set of measure zero with a graphon where every column maximizes $W$. In particular, every column must have the same worth.
@@
First we show that almost every column maximizes worth. Since we are working in $L^{2}$, where functions that differ on sets of measure zero are considered equivalent, we can get an equivalent representative by replacing any columns that don’t maximize worth with ones that do.
Let $W_{max}$ be the supremum of the worths of all possible columns. If there is a set of columns of positive measure whose worths are strictly less than $W_{max}$ then for some $\delta>0$ there is a set of columns of positive measure whose worths are all bounded above by $W_{max}-2\delta$. Replacing a subset $I_{s}$ of measure $s$ of such columns with a column $C$ whose worth is within $\delta$ of $W_{max}$ will increase $F$ by at least $s\delta+O(s^{2})$, contradicting Proposition 9.
(Note that “change a set of columns to CC” leads to an ambiguity for $g(x,y)$ when both $x$ and $y$ are in $I_{s}$. We can resolve this ambiguity by setting $g(x,y)=1$ on $I_{s}\times I_{s}$, or by setting $g(x,y)=0$, or by picking any other symmetric function in this square. Since the square where the ambiguity occurs has area $s^{2}$, different choices will yield values of $F$ that differ by $O(s^{2})$, which does not affect the violation of Proposition 9.) ∎
Having established the variational equations for nonsingular entropy maximizers, both for points with Theorem 7 and for columns with Theorem 11, we return to the (non)existence of singular entropy maximizers. If $g$ is a singular entropy maximizer, we call $\tau(g)$ a singular value of $t$ for the given edge density $e=\varepsilon(g)$. We claim that singular $t$’s are too sparse to matter. We begin with their measure.
Lemma 12.
For each $e\in(0,1)$, the set of singular $t$-values has measure zero.
Proof.
As explained below, the Boltzmann entropy function $\mathbb{B}(e,t)=\max_{g\in\mathcal{W}_{e,t}}S(g)$ for fixed $e$ is monotonically increasing in $t$ on $(t_{min},e^{3})$ and monotonically decreasing on $(e^{3},e^{3/2})$. This makes $d\mathbb{B}$ a finite measure on $(t_{min},e^{3})$ and makes $d\mathbb{B}$ a finite measure on $(e^{3},e^{3/2})$. We will show that the measure is singular at all singular $t$-values. The theorem then follows from the fact that the support of the singular part of a finite measure on an interval has zero Lebesgue measure.
To see the monotonicity of $\mathbb{B}$, let $g_{0}$ be an entropy maximizer at $(e,t)$ with $t<e^{3}$ and consider the family of graphons
(28) $g_{s}(x,y)=se+(1-s)g_{0}(x,y),$
which is defined for all $s\in[0,1]$. $S(g_{s})$ is an increasing function of $s$ (thanks to the concavity of $H(u)$) and in particular $S(g_{s})>S(g_{0})$ for all $s>0$. Since $\tau(g_{s})$ goes from $t$ to $e^{3}$ as $s$ goes from 0 to 1, and since $S(g_{s})$ is a lower bound for $\mathbb{B}(e,\tau(g_{s}))$, $\mathbb{B}(e,t^{\prime})>\mathbb{B}(e,t)$ for every $t^{\prime}\in(t,e^{3})$.
We also note that $\tau(g_{s})>t$ for all $s>0$, since otherwise, by the intermediate value theorem, there would be a positive $s$ with $\tau(g_{s})=t$. Since $S(g_{s})>S(g_{0})$, as we just discussed, that would contradict the optimality of $g_{0}$.
Now suppose that $t$ is singular, with a singular entropy maximizer $g_{0}$. Since $g_{0}(x,y)$ equals 0 or 1 on a set of positive measure, $S(g_{s})-S(g_{0})$ scales as $s\ln(1/s)$ as $s\to 0$. However, $\tau(g_{s})$ is a polynomial in $s$ and cannot grow faster than linearly for small $s$. Thus
(29) $\lim_{s\to 0^{+}}\frac{S(g_{s})-S(g_{0})}{\tau(g_{s})-\tau(g_{0})}=+\infty.$
Since $\mathbb{B}(e,t)=S(g_{0})$, and since $S(g_{s})$ is a lower bound for $\mathbb{B}(e,\tau(g_{s}))$, $\mathbb{B}$ must be increasing infinitely fast at $t$. That is, $t$ is a singular point of the measure $d\mathbb{B}$.
The exact same arguments work for $t>e^{3}$, only with $\tau(g_{s})$ being a decreasing function of $s$, with $\mathbb{B}$ being a decreasing function of $t$, and with $\mathbb{B}$ decreasing at infinite rate at singular $t$-values. ∎
We note that Lemma 12 proves there are no singular entropy maximizers in the regions studied in this paper.
Theorem 13.
Fix the edge density $e$ and consider an open interval $I$ of triangle densities $t$. Suppose that there is a differentiable function $f(t)$, defined for all $t\in I$, that equals the Boltzmann entropy $\mathbb{B}(e,t)$ whenever $t$ is non-singular. Then every $t\in I$ is non-singular.
Proof.
The Boltzmann entropy is never differentiable at $t=e^{3}$ [44], so we only need to consider intervals $I$ that are either above or below the Erdős-Rényi curve. By Lemma 12, the set of singular $t$-values in $I$ has measure zero, so the complement of that set is dense in $I$. Since $\mathbb{B}(e,t)$ is monotonic on $I$ and equals a (differentiable and therefore) continuous function $f(t)$ on a dense subset of $I$, it must equal $f(t)$ on all of $I$. But then $\mathbb{B}(e,t)$ is differentiable in $t$ for all $t\in I$, so there are no singular $t$-values in $I$.
Theorems 11 and 13 give us a strategy by which we can now prove (with much work!) Theorems 1, 2 and 4. In each case, we use variational equations that apply whenever $t$ is non-singular. For all such $t$, we show that the optimal graphon must take a certain form, with an entropy that is (the restriction of) a smooth function of $t$. Theorem 13 then implies that there are no singular $t$’s and that our calculations apply to all $t$. That is, while we cannot exclude singular entropy maximizers a priori, we are able to exclude them a posteriori.
We note that the analysis of optimizing the function $F$ in equation (20) is the starting point of our discussion of ERGMs in Section 6.